# Dynamic annihilation pathways of magnetic skyrmions

### v1

@@ Abstract
The investigation of magnetic solitons often relies on numerical modeling to determine key features such as stability, annihilation, nucleation, and motion. However, as soliton sizes approach atomic length scales, the accuracy of these predictions become increasingly sensitive to the details of the numerical model. Here, we study the annihilation of two-dimensional magnetic skyrmions using a pseudospectral approach and compare its performance to that of conventional micromagnetic simulations. A central distinction between the models lies in their treatment of the exchange interaction, which governs the magnon dispersion relation and plays a crucial role to balance the uniaxial anisotropy to stabilise skyrmions. We demonstrate that both the choice of model and spatial discretisation significantly influence the skyrmion dynamics and the magnetic field required for annihilation. The pseudospectral model provides a consistent description across length scales and captures complex behaviours such as skyrmion breathing on its path to annihilation. Our results have direct implications in the state-of-the-art modeling of skyrmions and other two-dimensional textures and will impact the modeling of three-dimensional textures such as hopfions.
@@
Nanoscale magnetic textures continue to attract intense research interest due to their rich underlying physics [1, 2] and promising technological applications. [3, 4, 5, 6, 7]. Even though many possible magnetic textures have been theoretically known for decades [1], only recent advances in fabrication and imaging techniques [8, 9, 10, 11, 12, 13, 14] have demonstrated their existence and opened pathways for their control. Topological textures [2] are of particular interest because their existence is protected by an energy barrier. Much effort has been devoted to magnetic skyrmions [3, 15, 16] which are topologically protected in two dimensions (2D), and, more recently, to magnetic hopfions [17, 18, 19, 20, 21, 22] which are topologically protected in three dimensions (3D). Despite their topological protection, it is challenging to preserve the stability of such textures. In the case of skyrmions, their topological protection is strictly in 2D [2], meaning that their stability in thin films is compromised by the spontaneous nucleation of hedgehog-antihedgehog pairs, or Bloch points [23, 24], that quickly annihilate at physical edges [25, 26]. In the case of hopfions, proximity to physical surfaces leads to their annihilation, although several experiments have provided evidence for their stabilization [19, 22].
@@
As material dimensions continue to shrink and topological textures approach atomic scales, the discrete nature of solid-state matter can no longer be neglected. Most theoretical investigations of topological textures rely on a continuum model of the magnetic exchange interaction [1, 27, 28, 29]. However, this approach introduces assumptions that break down at the atomic scale, e.g., a singular magnetisation vector at a Bloch point. One solution to this problem is to employ discrete numerical models such as atomistic spin dynamics (ASD) [30, 31]. Such models provide means to accurately describe Bloch points [32] and skyrmion annihilation [33, 34]. However, ASD simulations are constrained to small dimensions due to computational resource limitations. Another solution is to use micromagnetic simulations [35, 36, 37, 38, 39]. While this method models the exchange interaction using a continuum approximation, it is implemented numerically on a spatial grid of discrete cells. For this reason, micromagnetic simulations are assumed to be valid for cell dimensions larger than the exchange length and may become unphysical as cell dimensions approach the atomic lattice constant.
@@
ASD and micromagnetic simulations can complement each other, thus multiscale approaches have been sought [32, 40]. For example, smooth skyrmion motion was achieved in a lattice transitioning from an atomic to a continuum description [41]. Despite such successes, multiscale approaches consider a transition between different models which requires knowledge of the underlying geometry and a sufficiently large computational domain. Alternatively, a pseudospectral Landau-Lifshitz (PS-LL) description of magnetisation dynamics that models atomic and continuum dynamics on equal footing was introduced [42]. The PS-LL model leverages the fact that the magnon dispersion relation captures the relevant physical interactions at every spatial and temporal scale so that transition to larger cells is seamless in principle. This model has so far proven successful in describing ultrafast remagnetisation due to transient gratings [42, 43]. Here, we use the PS-LL to investigate the annihilation pathways of magnetic skyrmions.
@@
In this work, we demonstrate that the choice of numerical model and spatial discretisation can lead to quantitatively different skyrmion dynamics predictions. Using the PS-LL model, implemented with two distinct exchange interaction kernels, alongside results from the widely used micromagnetic simulation package MuMax3 [36], we examine the contraction and annihilation of a Néel-type skyrmion in a 2D magnetic thin film with uniaxial anisotropy and interfacial Dzyaloshinskii–Moriya interaction (DMI). The dynamics are driven by an increasing out-of-plane magnetic field. While all models show similar behaviour initially, striking differences emerge as the skyrmion radius approaches the exchange length. The PS-LL model accurately describes the dynamics at both the atomic and micromagnetic limits. However, micromagnetic models either overestimate or underestimate the exchange energy in the system, leading to an incorrect prediction of the skymrion’s topological protection and thus its eventual annihilation. Our work demonstrates that, even within the micromagnetic regime, a pseudospectral approach provides a much more accurate description of the magnetisation dynamics because the exchange interaction is described appropriately up to the first Brillouin zone. Such accuracy is essential for modeling the nucleation, stability, and annihilation of skyrmions as well as its manipulation by external stimuli. More broadly, these considerations are essential to correctly model three-dimensional solitons, including hopfions.
Results
Two-dimensional PS-LL model
Figure 1: Dispersion relations for different numerical models and spatial resolutions. Magnon dispersion relations extracted from dynamic simulations using three different models: the PS-LL model with the full magnon kernel (a and b), the PS-LL model with the micromagnetic kernel (c and d), and MuMax3 (e and f). The top row shows results at an atomic-scale discretization of 0.4 nm, while the bottom row uses a micromagnetic-scale discretization of 5.0 nm. a The magnon kernel at an atomic resolution reproduces the cosine-shaped dispersion relation for the first Brillouin zone. b The magnon kernel at 5.0 nm cell size produces a dispersion curve in a confined region (12.5 times smaller than panel a) of the first Brillioun zone, revealing a parabolic form consistent with the low-$k$ limit. c The micromagnetic kernel an an atomic resolution produces a parabolic dispersion which approximates this dispersion well near the origin but fails to capture short-wavelength behaviour. d The micromagnetic kernel at 5.0 nm produces a dispersion curve that shows excellent agreement with panel b, confirming the expected parabolic behaviour at this scale. e MuMax3 at an atomic resolution closely matches the magnon kernel at atomic resolution due to its nearest-neighbour finite-difference scheme. f MuMax3 at 5.0 nm produces a cosine-shaped dispersion due to the use of a finite-difference scheme and therefore deviates from the other two models at this scale.
@@ Results
where $\mathbf{m}$ denotes the magnetisation vector normalized by the saturation magnetisation, $M_{s}$, $\alpha$ is the Gilbert damping constant which can be used in the Landau-Lifshitz form is $\alpha\ll 1$, and the effective field field $\mathbf{H}_{\mathrm{eff}}$ is given by
@@
includes an external magnetic field with magnitude $H_{0}$ and a uniaxial anisotropy field with magnitufde $H_{k}$. The nonlocal contributions are incorporated in Fourier space through the term $\kappa(\mathbf{k})\mathbf{\hat{m}}$, where $\kappa(\mathbf{k})$ is a normalized nonlocal interaction kernel including exchange and Dzyaloshinskii-Moriya interaction (DMI), $\mathbf{\hat{m}}$ is the Fourier transform of the magnetisation, and $\mathcal{F}^{-1}$ represents the inverse Fourier transform.
For the exchange interaction in 2D, the contribution to the kernel is taken as the dispersion relation scaled by the saturation magnetisation [43],
@@
where $\lambda_{ex}$ is the exchange length and $a$ is the lattice constant. This dispersion relation accurately captures spin dynamics across the entire first Brillouin zone (FBZ). The dispersion relation reduces to the well known micromagnetic, long wavelength limit when $\mathbf{k}\rightarrow 0$
@@
Therefore, the PS-LL model allows for a smooth transition from the atomistic to the continuum limits based solely on the numerical discretisation [42].
The DMI contribution is considered to be interfacial, assuming a symmetry-breaking direction along the out-of-plane $\hat{\mathbf{z}}$ axis. Because DMI is asymmetric, it is convenient to represent the kernel as field components for each spatial direction independently:
@@
where $D$ is the interfacial DMI constant in units of J m-2.
@@
It is important to note that the DMI kernel is complex, which corresponds to a phase shift in Fourier space. Additionally, the presence of sine terms indicates that the kernel is an odd function, a property that, together with the phase shift, is the origin of nonreciprocity in the PS-LL description. Despite the kernel being complex, its inverse Fourier transform is real-valued, as expected from a field. In the long-wavelength limit ($\mathbf{k}\rightarrow 0$), a Taylor expansion of the sine functions shows that the kernel becomes linear in $k$, and the lattice constant vanishes. This is the regime where DMI is traditionally considered to be important as it produces a wavenumber shift in the dispersion relation [44]. However, as wavelengths become shorter, the DMI scaling in Eqs (6) can increase by a factor $1/a\approx 2.5$. The derivation of these DMI kernel components is provided in the Supplementary Material SI 1.
@@
The two energy contributions required to stabilize solitons [1] are contained within the effective field $\mathbf{H}_{\mathrm{eff}}$ from Eq. (2). The uniaxial anisotropy, contained in the local field term $\mathbf{h}_{l}$ introduces nonlinearities, while the nonlocal field term $\mathcal{F}^{-1}\left\{\kappa(\mathbf{k})\mathbf{\hat{m}}\right\}$ accounts for dispersion. Most theoretical understanding of magnetic solitons can be obtained within the micromagnetic limit [1, 29]. However, exact solutions are typically nonlinear, so their dynamics and decay can generally only be treated approximately [38, 39]. Numerical methods are therefore essential for accurately describing soliton behaviour in its entirety. For this reason, it is critical to assess the limitations that arise from the specific implementation of numerical models. In particular, since the magnon dispersion relation quantifies the energy of non-collinear spin configurations, a consistent dispersion relation across spatial scales is required for correct soliton modeling.
@@
To demonstrate grid independence and the effect of numerical approximations, Fig. 1 shows the dispersion relations extracted from simulations using the PS-LL model with the magnon kernel $\omega(\mathbf{k})$, the PS-LL model with the micromagentic kernel $\omega_{\mathrm{\mu}}(\mathbf{k})$, and MuMax3 [36], which is an open-source GPU-accelerated software that calculates the exchange interaction based on a finite difference method. For these simulations, we use material parameters consistent with Pt/Co/X multilayers, where X represents various metallic materials, $M_{s}=1120$ kA m-1, exchange stiffness of $A=10.0$ pJ m-1, uniaxial anisotropy field of $H_{k}=1370$ kA m-1, and DMI constant of $D=1.50$ mJ m-2 [45, 46]. The dispersion relation is obtained from the dynamic evolution of a Gaussian initial condition, as detailed in Methods. To isolate the effect of the dispersion kernel, we approximate the demagnetising field by setting a local contribution with magnitude equal to the saturation magnetisation, $M_{\mathrm{s}}$. This approximation captures the dominant out-of-plane demagnetisation effects in thin films, which is the regime considered here.
The exchange interaction dominates the shape of the dispersion relation. As shown in Fig. 1a and Fig. 1b, the PS-LL model incorporating the magnon kernel for the exchange interaction yields, as expected, the provided dispersion relation across all wavevectors and for each cell size. Specifically, Fig. 1a illustrates the dispersion obtained using an atomic cell size of 0.4 nm, while Fig. 1b shows the corresponding result using a micromagnetic cell size of 5 nm. The micromagnetic cell size returns a dispersion that closely follows a $k^{2}$ dependence in agreement with Eq. (5). This behaviour is expected near $k=0$, where the magnon dispersion is well approximated by the leading term of its Taylor expansion.
For comparison, Figs. 1c and 1d show the dispersion relations obtained using the micromagnetic kernel for atomic and micromagnetic cell sizes, respectively. In both cases, the result is purely parabolic, as expected. The close agreement between Fig. 1b and Fig. 1d further confirms the grid independence of the PS-LL model using the magnon kernel.
@@
We now compare these results with those obtained using MuMax3. Interestingly, for the atomic-scale discretisation shown in Fig. 1e, the dispersion closely matches that of the PS-LL model with the magnon kernel. This agreement arises from MuMax3’s finite-difference implementation of the exchange interaction, which uses a nearest-neighbour approximation. Such an approximation coincides with the Taylor expansion of a cosine function. This connection becomes even more apparent at the micromagnetic scale, shown in Fig. 1f, where the resulting dispersion also exhibits a cosine functional form. Consequently, the numerically computed dispersion relation depends on the manner in which the exchange interaction is implemented: whether through increasing the finite-difference stencil, as in OOMMF [37], including higher-order terms in the dispersion expansion [38], or adopting a finite-element approach.
@@
Discrepancies in the dispersion relation are often overlooked, as most magnetic textures span multiple cells, where the dispersion is approximately parabolic and consistent across models. However, as we demonstrate below, these discrepancies become relevant when examining the annihilation of topological textures such as skyrmions. Likewise, these discrepencies could also impact the nucleation and stability of such textures.
Skyrmion initial condition

### v3

@@ Abstract
The investigation of magnetic solitons often relies on numerical modeling to determine key features such as stability, annihilation, nucleation, and motion. However, as soliton sizes approach atomic length scales, the accuracy of these predictions becomes increasingly sensitive to the details of the numerical model. Here, we study the annihilation of two-dimensional magnetic skyrmions using a pseudospectral approach and compare its performance to that of conventional micromagnetic simulations. A central distinction between the models lies in their treatment of the exchange interaction, which governs the magnon dispersion relation and plays a crucial role in balancing the uniaxial anisotropy to stabilize skyrmions. We demonstrate that both the choice of model and the spatial discretization significantly influence skyrmion dynamics and the magnetic field required for annihilation. The pseudospectral model provides a consistent description across length scales and captures complex behaviors such as skyrmion breathing on the path toward annihilation. Our results have direct implications for the state-of-the-art modeling of skyrmions and other two-dimensional textures and will also impact the modeling of three-dimensional textures such as hopfions.
I Introduction
@@
Nanoscale magnetic textures continue to attract intense research interest due to their rich underlying physics [1, 2] and promising technological applications [3, 4, 5, 6, 7, 8, 9]. Even though many possible magnetic textures have been theoretically known for decades [1], only recent advances in fabrication and imaging techniques [10, 11, 12, 13, 14, 15, 16] have demonstrated their existence and opened pathways for their control. Topological textures [2] are of particular interest because their existence is protected by an energy barrier. Much effort has been devoted to magnetic skyrmions [3, 17, 18], which are topologically protected in two dimensions (2D), and, more recently, to magnetic hopfions [19, 20, 21, 22, 23, 24], which are topologically protected in three dimensions (3D). Despite their topological protection, it is challenging to preserve the stability of such textures. In the case of skyrmions, their topological protection is strictly in 2D [2], meaning that their stability in thin films is compromised by the spontaneous nucleation of hedgehog-antihedgehog pairs, or Bloch points [25, 26], which quickly annihilate at physical edges [27, 28]. In the case of hopfions, proximity to physical surfaces can lead to their annihilation, although several experiments [21, 24] and analytical studies [29] have provided evidence for their stabilization.
@@
As material dimensions continue to shrink and topological textures approach atomic scales, the discrete nature of solid-state matter can no longer be neglected. Most theoretical investigations of topological textures rely on a continuum model of the magnetic exchange interaction [1, 30, 31, 32]. However, this approach introduces assumptions that break down at the atomic scale, e.g., a singular magnetization vector at a Bloch point. One solution to this problem is to employ discrete numerical models such as atomistic spin dynamics (ASD) [33, 34]. Such models provide a means to accurately describe Bloch points [35] and skyrmion annihilation [36, 37]. However, ASD simulations are constrained to small system sizes due to computational resource limitations. Another solution is to use micromagnetic simulations [38, 39, 40, 41, 42]. This method models the exchange interaction using a continuum approximation, and therefore is strictly valid only for small wavevectors or large cell sizes. However, because the model is implemented numerically on a spatial grid of discrete cells, accurate solutions are typically achieved with small cell dimensions. For this reason, micromagnetic simulations are assumed to be valid for cell dimensions smaller than the exchange length but may nevertheless become unphysical as cell dimensions approach the atomic lattice constant.
@@
ASD and micromagnetic simulations can complement each other, thus multiscale approaches have been sought [35, 43]. For example, smooth skyrmion motion was achieved in a lattice transitioning from an atomic to a continuum description [44]. Despite such successes, multiscale approaches involve a transition between different models, which requires knowledge of the underlying geometry and a sufficiently large computational domain. Alternatively, one could use a pseudospectral Landau-Lifshitz (PS-LL) description of magnetization dynamics, which models atomic and continuum dynamics on equal footing [45]. The PS-LL model leverages the fact that the magnon dispersion relation captures the relevant physical interactions at every spatial and temporal scale, so that the transition to larger cells is seamless in principle. This model has so far proven successful in describing ultrafast remagnetization induced by transient gratings [45, 46]. Here, we extend the PS-LL to include the antisymmetric exchange interaction derived from first principles and apply it to investigate the annihilation pathways of magnetic skyrmions.
@@
In this work, we demonstrate that the choice of numerical model and spatial discretization can lead to quantitatively different predictions of skyrmion dynamics. Using the PS-LL model, implemented with two distinct exchange interaction kernels, alongside results from the widely used micromagnetic simulation package MuMax3 [39], we examine the contraction and annihilation of a Néel-type skyrmion in a 2D magnetic thin film with uniaxial anisotropy and interfacial Dzyaloshinskii–Moriya interaction (DMI). The dynamics are driven by an increasing out-of-plane magnetic field. While all models show similar behavior initially, striking differences emerge as the skyrmion radius approaches the exchange length. The PS-LL model accurately describes the dynamics at both the atomic and micromagnetic limits. However, micromagnetic models either overestimate or underestimate the exchange energy in the system, leading to an incorrect prediction of the skyrmion’s topological protection and thus its eventual annihilation. Our work demonstrates that, even within the micromagnetic regime, a pseudospectral approach provides a much more accurate description of the magnetization dynamics because the exchange interaction is described appropriately up to the first Brillouin zone. Such accuracy is essential for modeling the nucleation, stability, and annihilation of skyrmions as well as their manipulation by external stimuli. From a much broader point of view, our investigation highlights the limitations of standard solvers and computational tools that must be carefully considered as physical systems approach their ultimate scales and dynamics escape the standard models and approximations. Dispersion relations can be typically obtained by first principles and experiments, e.g., neutron scattering, for a variety of physical systems, and from there can be directly implemented in pseudospectral approaches.
II Two-dimensional PS-LL model
@@
where $\mathbf{m}$ denotes the magnetization vector normalized by the saturation magnetization, $M_{s}$, $\alpha$ is the Gilbert damping constant which can be used in the Landau-Lifshitz form provided $\alpha\ll 1$, and the effective field $\mathbf{H}_{\mathrm{eff}}$ is given by
@@
consists of an external magnetic field of magnitude $H_{0}$ and a uniaxial anisotropy field of magnitude $H_{k}$. The nonlocal contributions are incorporated in Fourier space through the term $\kappa(\mathbf{k})\mathbf{\hat{m}}$, where $\kappa(\mathbf{k})$ is a normalized nonlocal interaction kernel including exchange and Dzyaloshinskii-Moriya interaction (DMI), $\mathbf{\hat{m}}$ is the Fourier transform of the magnetization, and $\mathcal{F}^{-1}$ represents the inverse Fourier transform.
For the exchange interaction in 2D, the contribution to the kernel is taken as the dispersion relation scaled by the saturation magnetization [46],
@@
where $\lambda_{ex}$ is the exchange length and $a$ is the lattice constant. This dispersion relation can be obtained from first principles by considering the nearest-neighbor interaction between localized moments, i.e., under the assumption of a Heisenberg exchange model [47]. Higher-order exchange terms could be incorporated to represent next-nearest-neighbor interactions [29], but we restrict our current investigation to nearest-neighbor interactions only. Under this assumption, Eq. (4) provides an accurate representation of the spin dynamics across the entire first Brillouin zone (FBZ). The dispersion relation reduces to the well-known micromagnetic, long-wavelength limit when $\mathbf{k}\rightarrow 0$
@@
Therefore, the PS-LL model allows for a smooth transition from the atomistic to the continuum limits, based solely on the numerical discretization [45].
The DMI is considered to be interfacial, with a symmetry-breaking direction along the out-of-plane $\hat{\mathbf{z}}$ axis. Its contribution to the dispersion relation can be derived from the DMI energy, considering a square lattice, as detailed in Appendix A. Because DMI is asymmetric, it is convenient to represent the kernel as field components for each spatial direction independently:
@@
where $D$ is the interfacial DMI constant, with units of J m-2.
@@
It is important to note that the DMI kernel is complex, which corresponds to a phase shift in Fourier space. Additionally, the presence of sine terms indicates that the kernel is an odd function, a property that, together with the phase shift, is the origin of nonreciprocity in the PS-LL description. Although the kernel is complex, its inverse Fourier transform is real-valued, as expected from a physical field. In the long-wavelength limit ($\mathbf{k}\rightarrow 0$), a Taylor expansion of the sine functions shows that the kernel becomes linear in $k$, and the lattice constant vanishes. This is the regime where DMI is traditionally considered to be important as it produces a wavenumber shift in the dispersion relation [48]. However, as wavelengths become shorter, the DMI scaling in Eqs (6) can increase by a factor $1/a\approx 2.5$.
Figure 1: Magnon dispersion relations extracted from dynamic simulations using three different models: the PS-LL model with the full magnon kernel, (a) and (b), the PS-LL model with the micromagnetic kernel, (c) and (d), and MuMax3 (e) and (f). The top row shows results at an atomic-scale discretization of 0.4 nm, while the bottom row uses a micromagnetic-scale discretization of 5.0 nm. (a) The magnon kernel at an atomic resolution reproduces the cosine-shaped dispersion relation for the first Brillouin zone. (b) The magnon kernel at 5.0 nm cell size produces a dispersion curve in a confined region [12.5 times smaller than panel (a)] of the first Brillouin zone, revealing a parabolic form consistent with the low-$k$ limit. (c) The micromagnetic kernel an atomic resolution produces a parabolic dispersion which approximates this dispersion well near the origin but fails to capture short-wavelength behavior. (d) The micromagnetic kernel at 5.0 nm produces a dispersion curve that shows excellent agreement with panel (b), confirming the expected parabolic behavior at this scale. (e) MuMax3 at an atomic resolution closely matches the magnon kernel at atomic resolution due to its nearest-neighbor finite-difference scheme. (f) MuMax3 at 5.0 nm produces a cosine-shaped dispersion due to the use of a finite-difference scheme and therefore deviates from the other two models at this scale. The color scheme is found in Ref. [49].
@@
The two energy contributions required to stabilize solitons [1] are contained within the effective field $\mathbf{H}_{\mathrm{eff}}$ from Eq. (2). The uniaxial anisotropy, contained in the local field term $\mathbf{h}_{l}$ introduces nonlinearities, while the nonlocal field term $\mathcal{F}^{-1}\left\{\kappa(\mathbf{k})\mathbf{\hat{m}}\right\}$ accounts for dispersion. Most theoretical understanding of magnetic solitons can be obtained within the micromagnetic limit [1, 32]. However, exact solutions are typically nonlinear, so their dynamics and decay can generally only be treated approximately [41, 42]. Numerical methods are therefore essential for accurately describing soliton behavior in its entirety. For this reason, it is critical to assess the limitations that arise when implementing a specific numerical model. In particular, since the magnon dispersion relation quantifies the energy of non-collinear spin configurations, a consistent dispersion relation across spatial scales is required for correct soliton modeling. It is also important to note that the dispersion relation is derived under the assumption of small-amplitude, linear waves, but its physical origin does not rely on linearity. This is the key advantage of the pseudospectral model: the dispersion relation serves as a kernel that quantifies the Heisenberg interaction and can, in fact, be derived directly from first principles as shown above for the DMI and as traditionally done to obtain the magnon dispersion relation [47].
@@
To demonstrate grid independence and the effect of numerical approximations, Fig. 1 shows the dispersion relations extracted from simulations using the PS-LL model with the magnon kernel $\omega(\mathbf{k})$, the PS-LL model with the micromagnetic kernel $\omega_{\mathrm{\mu}}(\mathbf{k})$, and MuMax3 [39], which is an open-source GPU-accelerated software that calculates the exchange interaction based on a finite difference method. For these simulations, we use material parameters consistent with Pt/Co/X multilayers, where X represents various metallic materials, $M_{s}=1120$ kA m-1, exchange stiffness $A=10.0$ pJ m-1, uniaxial anisotropy field $H_{k}=1370$ kA m-1, and DMI constant $D=1.50$ mJ m-2 [50, 51]. For these material parameters, the exchange length is $\lambda_{\mathrm{ex}}=\sqrt{2A/(\mu_{0}M_{s}^{2})}=3.56$ nm.
The dispersion relation is obtained from the dynamic evolution of a Gaussian initial condition set with a standard deviation equal to the cell size, effectively exciting the full available spectrum due to its broadband nature. By recording the decay of this initial condition, the magnon modes across the entire first Brillouin zone are populated. A fast Fourier transform is then applied to the resulting time-dependent $m_{z}$ magnetization component in both space and time from which a two-dimensional cross-section yields the dispersion relations shown in Fig. 1. A Gilbert damping of $\alpha=0.005$ is used to maximize the contrast in Fourier space.
Because the material has perpendicular magnetic anisotropy (PMA), the Gaussian initial condition was set as a deviation from $m_{z}=1$, with amplitude 0.9, such that the minimum $m_{z}=0.1$ occurred at the Gaussian’s center. The condition $|\mathbf{m}|=1$ was enforced by setting $m_{x}=\sqrt{1-m_{z}^{2}}$ and $m_{y}=0$ as the initial condition. We note that the Gaussian was set with an amplitude smaller than 1 to avoid the onset of nonlinear effects that would lead to the mixing of spectral components. Therefore, the dispersion relation can be reproduced for any Gaussian amplitude under 1.
To isolate the effect of the dispersion kernel, the demagnetizing field was approximated by setting a local contribution with magnitude equal to the saturation magnetization, $M_{\mathrm{s}}$. In other words, the anisotropy field $H_{k}\rightarrow H_{k}-M_{s}$. This approximation captures the dominant out-of-plane demagnetization effects in thin films, which is the regime considered here.
[…]
@@
We now compare these results with those obtained using MuMax3. Interestingly, for the atomic-scale discretization shown in Fig. 1(e), the dispersion closely matches that of the PS-LL model with the magnon kernel. This agreement arises from MuMax3’s finite-difference implementation of the exchange interaction, which uses a nearest-neighbor approximation. Such an approximation coincides with the Taylor expansion of a cosine function. This connection becomes even more apparent at the $5.0$ nm cell-size, shown in Fig. 1(f), where the resulting dispersion also exhibits a cosine functional form. It is important to stress that even though the cell-size is larger than the exchange length, the simulation is well-behaved and numerical instabilities are not observed at any time. From these results, we conclude that the numerically computed dispersion relation depends on the manner in which the exchange interaction is implemented: whether through increasing the finite-difference stencil, as in OOMMF [40], including higher-order terms in the dispersion expansion [41], or adopting a finite-element approach.
@@
As a consequence of the linearization of the Landau-Lifshitz equation, the dispersion relation is often overlooked when considering nonlinear structures such as solitons. However, the dispersion relation encodes the coupling between space and time in the system and thus fully describes the interactions between magnetic moments, regardless of the system’s linearity. As we demonstrate below, the discrepancies between the dispersion relations from different models become relevant when examining the annihilation of topological textures such as skyrmions. Likewise, these discrepancies could also impact the nucleation and stability of such textures.
III Dynamic skyrmion annihilation